# Semantic MIMO Systems for Speech-to-Text Transmission

### v1

@@ IV-C MU-MIMO Experiments
In the MU-MIMO communication scenarios, the WER scores and the sentence similarity scores are calculated by averaging the scores of $K$ users. The performance comparison amongst the DeepSC-ST, DeepSC-ST 2.0, and the proposed SAC-ST is shown in Fig. 9. From the figure, the SAC-ST shows the superiority of providing accurate sentences to $K$ users under the tested channel environments.
The results of SAC-ST in MU-MIMO transmission systems with different antenna combinations and number of users are shown in Fig. 10. From the figure, when serving the same number of users, the semantic features passing through the MIMO channels with more antennas experience reduced channel impairment in the low SNR regime according to the distribution of equivalent SNR values $\widetilde{\boldsymbol{\sigma}}$ in Table III. However, as the number of users increases, the accuracy of the sentences produced for $K$ users diminishes because the limited essential information of each user is transmitted over the good SISO channels.
Fig. 11 represents the SAC-ST with various channel estimation approaches in the MU-MIMO transmission system. From the figures, the SAC-ST with the ChanEst network outperforms all the benchmarks, and it obtains approximate WER scores and sentence similarity scores compared to the SAC-ST with perfect CSI, which verifies the adaptability of the developed ChanEst network in both SU-MIMO and MU-MIMO communication systems.
@@
V Conclusions
In this article, we developed a semantic-aware communication system for speech-to-text transmission in single-user and multi-user MIMO communication systems, named SAC-ST. The transformer-enabled semantic communication system is first developed to achieve the speech-to-text task at the receiver user. A novel semantic-aware network is also proposed to recognize the critical semantic features and produce the importance matrix, which prompts the information exchange with high semantic fidelity by recovering the essential semantic information accurately. Furthermore, a neural network-enabled channel estimation network, named ChanEst network, is designed to boost the SAC-ST transmission over realistic channel environments. Simulation results verified the superiority of the proposed SAC-ST, especially in the low SNR regime, and proved the reliability of ChanEst network for MIMO channel estimation. Therefore, our proposed SAC-ST is a promising candidate for speech-to-text transmission over practical MIMO communication systems.

### v2

@@ IV-C MU-MIMO Experiments
In the MU-MIMO communication scenarios, the WER scores and the sentence similarity scores are calculated by averaging the scores of $K$ users. The performance comparison amongst the benchmarks, DeepSC-ST 2.0, and the proposed SAC-ST is shown in Fig. 10. From the figure, the SAC-ST shows the superiority of providing accurate sentences to $K$ users under the tested channel environments.
@@
The results of SAC-ST in MU-MIMO transmission systems with different antenna combinations and number of users are shown in Fig. 11. From the figure, when serving the same number of users, the semantic features passing through the MIMO channels with more antennas experience reduced channel impairment in the low SNR regime according to the distribution of equivalent SNR values $\widetilde{\boldsymbol{\sigma}}$ in Table IV. However, as the number of users increases, the accuracy of the sentences produced for $K$ users diminishes because the limited essential information of each user is transmitted over the good SISO subchannels.
@@
Fig. 12 illustrates the SAC-ST with various channel estimation approaches in the MU-MIMO transmission system. From the figures, the SAC-ST with the ChanEst network outperforms all the benchmarks, and it obtains approximate WER scores and sentence similarity scores compared to the SAC-ST with perfect CSI, which verifies the adaptability of the developed ChanEst network in both SU-MIMO and MU-MIMO communication systems.
Furthermore, the average model inference time of different systems under the SU-MIMO and MU-MIMO communication scenarios is provided in Table VI.
TABLE VI: Model inference time of different systems.
Inference Time (Second)
SU-MIMO SAC-ST w/ Perfect CSI 0.039
SAC-ST w/ ChanEst 0.047
MU-MIMO SAC-ST w/ Perfect CSI 0.086
SAC-ST w/ ChanEst 0.089
V Conclusions and Future Directions
In this article, we developed a semantic-aware communication system for speech-to-text transmission in single-user and multi-user MIMO communication systems, named SAC-ST. The transformer-enabled semantic communication system is first developed to achieve the speech-to-text task at the receiver. A novel semantic-aware network is also proposed to recognize the critical semantic features and produce the importance matrix, which prompts the information exchange with high semantic fidelity by recovering the essential semantic information accurately. Furthermore, a neural network-enabled channel estimation network, named ChanEst network, is designed to boost the SAC-ST transmission over realistic channel environments. Simulation results verified the superiority of the proposed SAC-ST, especially in the low SNR regime, and proved the reliability of ChanEst network for MIMO channel estimation. Therefore, our proposed SAC-ST is a promising candidate for speech-to-text transmission over practical MIMO communication systems.
In our work, the proposed semantic communication architecture is tailored for English. However, the language diversity in practical speech transmission scenarios imposes great challenges to the robustness of semantic communications, necessitating a comprehensive model to serve users with different language backgrounds, including dialects and accents. Therefore, the further research involves developing a robust semantic communication paradigm for multilingual environments.